# On the Stability of Electromechanical Switching Devices

### v3

@@ I Introduction
These switching devices are all based on a small single-coil reluctance actuator with a limited range of motion. In essence, a reluctance actuator is an electromagnet, constructed by wrapping a wire around a ferromagnetic core, together with a movable component known as armature (see Fig. 1). When the coil is supplied with power, the magnetic flux that flows through the core creates a magnetic force that pulls the armature toward the yoke. This force is always attractive, so the reverse motion is usually produced by a spring or simply by gravity. More advanced actuators can be designed from this concept, e.g., by including permanent magnets to reduce power consumption [5, 6] or a second coil to increase control possibilities [7, 8].
@@
Although there is much research on modeling, control, and estimation of solenoid actuators, on/off valves, and electromagnetic relays, a significant gap exists in the analysis area. In this paper, we present for the first time in the literature a stability analysis for electromechanical switching devices based on a single-coil reluctance actuator. The main contribution of the work is to explain why these systems exhibit hysteretic bistable behavior with respect to the supply voltage, i.e., why they are only stable at two given positions and why the switching in one direction occurs at a different voltage level than in the other direction. The study leverages a hybrid dynamical model, based on the insights and methodologies established in our previous research [13][15], to obtain a series of analytic expressions relating the physical parameters of the system to the switching conditions. In order to improve the understanding, the theoretical results are presented alongside plots generated using parameter values that may be representative of a typical electromechanical switching device. The results are discussed in depth, emphasizing the implications in the design of new switching devices and other possible practical uses. To validate the theoretical findings, we also present experimental results obtained from a real device, confirming the predicted hysteretic bistable behavior. Finally, to further demonstrate the practical value of the research, we describe and present the results of a parameter estimation procedure based on the derived expressions.
@@ II Modeling
The dynamics of the actuator is explained with the help of the diagram in Fig. 1, which contains all the information required to describe the electromagnetic dynamics of any single-coil actuator. A coil of $N$ turns is wrapped around the core, with an electric current $i$ flowing through it. This current, whose dynamics is governed by the electrical power supply circuit, is responsible for inducing the magnetic flux $\phi$ in the core. Due to the high permeability of the iron, the magnetic flux remains confined almost entirely in the core, with the exception of the segment where it crosses the air gap. As it will be seen later, the magnetic flux is directly related to the magnetic force, which always acts by attracting both ends of the air gap. The length of the air gap, denoted by $z$ in the figure, will be used as the variable that defines the position of the actuator. Regardless of the actuator design, the air gap length is always greater than or equal to zero. In addition to this constraint—negative air gaps have no physical meaning—all actuators include some type of mechanical constraints that limit the armature stroke. Without loss of generality, it can thus be stated that $z\in\left[z_{\mathrm{1}},\,z_{\mathrm{2}}\right]$, with $0\leq z_{\mathrm{1}}<z_{\mathrm{2}}$.
@@ II-A Free motion dynamics
One of the benefits of using a reluctance-based model is that this variable is also directly related to the magnetic—or reluctance—force. Given a reluctance $\mathcal{R}(z,\phi)$, this force is given [13] by
@@
The magnetic force is the only external excitation of the mass-spring-damper system that forms the actuator mechanism. In this work the spring will be considered to be linear and the damping force will be defined as a function $f_{\mathrm{d}}(\dot{z})$ under the condition $f_{\mathrm{d}}(0)=0$, which encompasses a wide range of frictional damping models with varying degrees of sophistication [30]. The dynamics of the motion, which is given by Newton’s second law, is therefore as follows,
$m\,\ddot{z}=F_{\mathcal{R}}(z,\phi)-k_{\mathrm{s}}\left(z-z_{\mathrm{s}}\right)-f_{\mathrm{d}}(\dot{z}),$ (7)
@@
$\dot{v}=f_{v}({x})=\frac{F_{\mathcal{R}}(z,\phi)}{m}-\frac{k_{\mathrm{s}}}{m}\left(z-z_{\mathrm{s}}\right)-\frac{1}{m}\,f_{d}(\dot{z}),$ (9)
@@ II-B Dynamics of switching actuators
Although there are some reluctance actuators designed to operate in continuous mode [31], all low-cost switching devices based on this technology are intended to switch between the two limit positions. This category includes electromechanical relays, whose purpose is to open and close electrical connections, and solenoid actuators, which are widely used as on/off valves in hydraulic and pneumatic circuits. The mechanical limits of these devices not only constrain the movement of the armature to a certain range, but also result in the occurrence of strong impacts at the end of the displacements. Furthermore, these impacts often lead to armature bouncing. Since the velocity changes instantaneously on impacts, this behavior can be considered as a case of hybrid dynamics.
@@ III Analysis in continuous mode
In this and the following section we analyze the equilibrium points and the stability of the system as functions of the supply voltage. As it will be seen, the variation of $u$ produces certain changes in the phase space that explain to a large extent the behavior of this class of devices. We present, whenever possible, the closed-form expressions of the points of interest as a function of the parameters. However, specific parameter values have been used to plot the results in the figures. These values, which may be representative of a typical solenoid actuator or electromagnetic relay, are shown in Table I.
@@
[…]
$\phi_{\mathrm{sat}}$ $20$ $\upmu$Wb $c$ $0.1$ Ns/m
@@ III-A Basic model
Figure 4: Basic model under continuous operation. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are represented with solid/dashed lines.
@@ III-B Model with magnetic saturation
The tangential bifurcations that exist in the basic model are also present in this case. These occur for those values of $u$ at which the curves $f_{\phi}(x,u)=0$ and $f_{v}(x)=0$ become tangent. There are two solutions: one for positive supply voltage $u=\tilde{u}_{\mathrm{b}}$ at $(\phi,z)=(\tilde{\phi}_{\mathrm{b}},\tilde{z}_{\mathrm{b}})$, and another one for negative voltage $u=-\tilde{u}_{\mathrm{b}}$ at $(\phi,z)=(-\tilde{\phi}_{\mathrm{b}},\tilde{z}_{\mathrm{b}})$. In this case it is also possible to obtain closed-form solutions for these points, but the resulting expressions are of such length that their practical use is limited. In fact, we do not present them in the manuscript for the sake of brevity. Comparing the curves in Figs. 3 and 5, however, it can be seen that the bifurcations in this model must necessarily occur at points farther away from the vertex of the parabola. That is, $\tilde{z}_{\mathrm{b}}<{z}_{\mathrm{b}}$ and $\tilde{\phi}_{\mathrm{b}}>{\phi}_{\mathrm{b}}$. The relation between $\tilde{u}_{\mathrm{b}}$ and ${u}_{\mathrm{b}}$, which can be obtained by solving $f_{\phi}(x,u)=0$ for each of the models, is given by
@@
Figure 6: Model with saturation under continuous operation. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are represented with solid/dashed lines.
@@ IV Analysis under hybrid dynamics
It is not difficult to realize that this analysis is much richer than in the continuous case, as the behavior of the system may change considerably depending on the values of $z_{\mathrm{1}}$ and $z_{\mathrm{2}}$. In this work we assume that $z_{\mathrm{1}}\in\left[0,{z}_{\mathrm{b}}\right)$—or $z_{\mathrm{1}}\in\left[0,\tilde{z}_{\mathrm{b}}\right)$, when considering saturation—a condition that is met in virtually all commercial devices. In the following we analyze, for the two models previously proposed, the different cases that may arise depending on the value of $z_{\mathrm{2}}$.
@@ IV-A Basic model
Figure 7: Basic model under hybrid dynamics. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are represented with solid/dashed lines. (a) Case 1: $z_{\mathrm{2}}>z_{\mathrm{s}}$. (b) Case 2: $z_{\mathrm{b}}<z_{\mathrm{2}}<z_{\mathrm{s}}$. (c) Case 3: $z_{\mathrm{1}}<z_{\mathrm{2}}<z_{\mathrm{b}}$.
@@ IV-B Model with magnetic saturation
Figure 8: Model with saturation under hybrid dynamics. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are represented with solid/dashed lines. (a) Case 1: $z_{\mathrm{2}}>z_{\mathrm{s}}$. (b) Case 2: $\tilde{z}_{\mathrm{b}}<z_{\mathrm{2}}<z_{\mathrm{s}}$. (c) Case 3: $z_{\mathrm{1}}<z_{\mathrm{2}}<\tilde{z}_{\mathrm{b}}$.
@@ V Experimental demonstration
The device used in the tests is a commercial single-pole double-throw (SPDT) electromechanical relay (see Fig. 10a). The experimental setup (see Fig. 10b) consists of a personal computer, a USB oscilloscope with a signal generator (PicoScope 4824), and an oscilloscope-relay link circuit. The personal computer, together with the oscilloscope, controls the applied signal and records the actual voltage, the relay excitation current and the status (open or closed) of the electrical contacts. The oscilloscope-relay link circuit (see Fig. 10c) consists of a non-inverting amplifier stage, a current shunt, and a resistive circuit to monitor the contacts status.
@@
Figure 10: Experimental setup. (a) Tested electromechanical relay without its enclosure. (b) Picture of the experimental setup. (c) Schematic diagram of the oscilloscope-relay link circuit.
Figure 10: Experimental setup. (a) Tested electromechanical relay without its enclosure. (b) Picture of the experimental setup. (c) Schematic diagram of the oscilloscope-relay link circuit.
@@
TABLE III: Estimated parameter values.
@@
[…]
@@ VI Conclusions
This paper has presented a comprehensive stability analysis for electromechanical switching devices, a category encompassing electromagnetic relays and solenoid actuators. One of the key objectives was to derive expressions for the voltage and magnetic flux at which these devices switch. To do this, we have leveraged two variants of a state-of-the-art hybrid dynamical model, including magnetic saturation in one of them for enhanced accuracy at the cost of greater complexity. Both models assume a one-to-one relationship between current and magnetic flux. This simplification yields valuable analytical expressions for initial understanding, rapid analysis, and design optimization. For applications involving high-coercivity materials or high-frequency scenarios, a more comprehensive analysis incorporating magnetic hysteresis may be beneficial, although this would probably imply numerical procedures.

### v4

@@ I Introduction
These switching devices are all based on a small single-coil reluctance actuator with a limited range of motion. In essence, a reluctance actuator is an electromagnet, constructed by wrapping a wire around a ferromagnetic core, together with a movable component known as armature (see Fig. 1). When the coil is supplied with power, a magnetic force that pulls the armature toward the yoke is created. This force is always attractive, so the reverse motion is usually produced by a spring or simply by gravity. More advanced actuators can be designed from this concept, e.g., by including permanent magnets to reduce power consumption [5, 6] or a second coil to increase control possibilities [7, 8].
@@
Although there is much research on modeling, control, and estimation of solenoid actuators, on/off valves, and electromagnetic relays, a significant gap exists in the analysis area. In this paper, we present for the first time in the literature a stability analysis for electromechanical switching devices based on a single-coil reluctance actuator. The main contribution of the work is to explain why these systems exhibit hysteretic bistable behavior with respect to the supply voltage, i.e., why they are only stable at two given positions and why the switching in one direction occurs at a different voltage level than in the other direction. The study leverages a hybrid dynamical model, based on the insights and methodologies established in our previous research [13][15], to obtain a series of analytic expressions relating the physical parameters of the system to the switching conditions. In order to improve the understanding of the theoretical results, the expressions obtained are presented alongside plots generated using parameter values that can be considered representative of a typical electromechanical switching device. The results are discussed in depth, emphasizing the implications in the design of new switching devices and other possible practical uses. To validate the theoretical findings, we also present experimental results obtained from a real device, confirming the predicted hysteretic bistable behavior. Finally, a parameter estimation procedure based on the derived expressions demonstrates the practical value of the research.
@@ II Modeling
The dynamics of the actuator is explained with the help of the diagram in Fig. 1. A coil of $N$ turns is wrapped around the core, with an electric current $i$ flowing through it. This current, whose dynamics is governed by the electrical power supply circuit, is responsible for inducing the magnetic flux $\phi$ in the core. Due to the high permeability of the iron, the magnetic flux remains confined almost entirely in the core, with the exception of the segment where it crosses the air gap. As it will be seen later, the magnetic flux is directly related to the magnetic force, which always acts by attracting both ends of the air gap. The length of the air gap, denoted by $z$ in the figure, will be used as the variable that defines the position of the actuator. Regardless of the actuator design, the air gap length is always greater than or equal to zero. In addition to this constraint—negative air gaps have no physical meaning—all actuators include some type of mechanical constraints that limit the armature stroke. Without loss of generality, it can thus be stated that $z\in\left[z_{\mathrm{1}},\,z_{\mathrm{2}}\right]$, with $0\leq z_{\mathrm{1}}<z_{\mathrm{2}}$.
@@ II-A Free motion dynamics
One of the benefits of using a reluctance-based model is that this variable is also directly related to the magnetic force. Given a reluctance $\mathcal{R}(z,\phi)$, this force is given [13] by
@@
The magnetic force is the only external excitation of the mass-spring-damper system that forms the actuator mechanism. In this work, we assume a linear spring and a damping force described by a function of the form $f_{\mathrm{d}}:\dot{z}\mapsto f_{\mathrm{d}}(\dot{z})$, with $f_{\mathrm{d}}(0)=0$. This latter assumption encompasses a wide range of frictional damping models with varying degrees of sophistication [30], which implies that the analyses in the following sections are more generalizable. Notably, any model that meets this assumption exhibits identical behavior at equilibrium. The dynamics of the motion, given by Newton’s second law, is therefore as follows,
$m\,\ddot{z}=-\frac{1}{2}\,k_{\mathcal{R}}\,\phi^{2}-k_{\mathrm{s}}\left(z-z_{\mathrm{s}}\right)-f_{\mathrm{d}}(\dot{z}),$ (7)
@@
$\dot{v}=f_{v}({x})=-\frac{k_{\mathcal{R}}}{2\,m}\,\phi^{2}-\frac{k_{\mathrm{s}}}{m}\left(z-z_{\mathrm{s}}\right)-\frac{1}{m}\,f_{\mathrm{d}}(\dot{z}),$ (9)
@@ II-B Dynamics of switching actuators
Although there are some reluctance actuators designed to operate in continuous mode [31], all low-cost switching devices based on this technology are intended to switch between the two limit positions. This category includes electromagnetic relays, whose purpose is to open and close electrical connections, and solenoid actuators, which are widely used as on/off valves in hydraulic and pneumatic circuits. The mechanical limits of these devices not only constrain the movement of the armature to a certain range, but also result in the occurrence of strong impacts at the end of the displacements. Furthermore, these impacts often lead to armature bouncing. Since the velocity changes instantaneously on impacts, this behavior can be considered as a case of hybrid dynamics.
@@ III Analysis in continuous mode
In this and the following section, we analyze the equilibrium points and the stability of the system as functions of the supply voltage. As it will be seen, the variation of $u$ produces certain changes in the phase space that explain to a large extent the behavior of this class of devices. We present, whenever possible, the closed-form expressions of the points of interest as a function of the parameters. The figures, included to improve the understanding of the theoretical results, have been generated using the parameters in Table I, which can be considered representative of many commercial solenoid actuators.
@@
$\mathcal{R}_{c0}$ $1.5\!\cdot\!10^{7}$ H-1 $m$ $1$ g
$\mathcal{R}_{g0}$ 00 H-1 $k_{\mathrm{s}}$ $55$ N/m
$k_{\mathcal{R}}$ $2\!\cdot\!10^{10}$ H-1/m $z_{\mathrm{s}}$ $15$ mm
$\phi_{\mathrm{sat}}$ $20$ $\upmu$Wb - -
@@ III-A Basic model
Figure 4: Basic model under continuous operation. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are plotted with solid/dashed lines.
@@ III-B Model with magnetic saturation
The tangential bifurcations that exist in the basic model are also present in this case. These occur for those values of $u$ at which the curves $f_{\phi}(x,u)=0$ and $f_{v}(x)=0$ become tangent. There are two solutions: one for positive supply voltage $u=\tilde{u}_{\mathrm{b}}$ at $(\phi,z)=(\tilde{\phi}_{\mathrm{b}},\tilde{z}_{\mathrm{b}})$, and another one for negative voltage $u=-\tilde{u}_{\mathrm{b}}$ at $(\phi,z)=(-\tilde{\phi}_{\mathrm{b}},\tilde{z}_{\mathrm{b}})$. In this case, it is also possible to obtain closed-form solutions for these points, but the resulting expressions are of such length that their practical use is limited. In fact, we do not present them in the manuscript for the sake of brevity. Comparing the curves in Figs. 3 and 5, however, it can be seen that the bifurcations in this model must necessarily occur at points farther away from the vertex of the parabola. That is, $\tilde{z}_{\mathrm{b}}<{z}_{\mathrm{b}}$ and $\tilde{\phi}_{\mathrm{b}}>{\phi}_{\mathrm{b}}$. The relation between $\tilde{u}_{\mathrm{b}}$ and ${u}_{\mathrm{b}}$, which can be obtained by solving $f_{\phi}(x,u)=0$ for each of the models, is given by
@@
Figure 6: Model with saturation under continuous operation. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are plotted with solid/dashed lines.
@@ IV Analysis under hybrid dynamics
It is not difficult to realize that this analysis is much richer than in the continuous case, as the behavior of the system may change considerably depending on the values of $z_{\mathrm{1}}$ and $z_{\mathrm{2}}$. In this work, we assume that $z_{\mathrm{1}}\in\left[0,{z}_{\mathrm{b}}\right)$—or $z_{\mathrm{1}}\in\left[0,\tilde{z}_{\mathrm{b}}\right)$, when considering saturation—a condition that is met in virtually all commercial devices. In the following, we analyze, for the two models previously proposed, the different cases that may arise depending on the value of $z_{\mathrm{2}}$.
@@ IV-A Basic model
Figure 7: Basic model under hybrid dynamics. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are plotted with solid/dashed lines. (a) Case 1: $z_{\mathrm{2}}>z_{\mathrm{s}}$. (b) Case 2: $z_{\mathrm{b}}<z_{\mathrm{2}}<z_{\mathrm{s}}$. (c) Case 3: $z_{\mathrm{1}}<z_{\mathrm{2}}<z_{\mathrm{b}}$.
@@ IV-B Model with magnetic saturation
Figure 8: Model with saturation under hybrid dynamics. Bifurcation diagram of equilibria in terms of $u$. Position (top) and magnetic flux (bottom). Stable/unstable equilibria are plotted with solid/dashed lines. (a) Case 1: $z_{\mathrm{2}}>z_{\mathrm{s}}$. (b) Case 2: $\tilde{z}_{\mathrm{b}}<z_{\mathrm{2}}<z_{\mathrm{s}}$. (c) Case 3: $z_{\mathrm{1}}<z_{\mathrm{2}}<\tilde{z}_{\mathrm{b}}$.
@@ V Experimental demonstration
The device used in the tests is a commercial single-pole double-throw (SPDT) electromagnetic relay (see Fig. 10a). The experimental setup (see Fig. 10b) consists of a personal computer, a USB oscilloscope with a signal generator (PicoScope 4824), and an oscilloscope-relay link circuit. The personal computer, together with the oscilloscope, controls the applied signal and records the actual voltage, the relay excitation current and the status (open or closed) of the electrical contacts. The oscilloscope-relay link circuit (see Fig. 10c) consists of a non-inverting amplifier stage, a current shunt, and a resistive circuit to monitor the contacts status.
@@
Figure 10: Experimental setup. (a) Tested electromagnetic relay without its enclosure. (b) Picture of the experimental setup. (c) Schematic diagram of the oscilloscope-relay link circuit.
@@
TABLE III: Estimated parameter values for the electromagnetic relay.
@@
$\mathcal{R}_{c0}$ $1.15\!\cdot\!10^{4}$ H-1 $\phi_{\mathrm{sat}}$ $148$ $\upmu$Wb
$\mathcal{R}_{g0}$ $4.49\!\cdot\!10^{4}$ H-1 $k_{\mathrm{s}}$ $94.65$ N/m
$k_{\mathcal{R}}$ $7.95\!\cdot\!10^{7}$ H-1/m $z_{\mathrm{s}}$ $2.85$ mm
@@ VI Conclusions
In this paper, we have presented a comprehensive stability analysis for electromechanical switching devices, a category encompassing electromagnetic relays and solenoid actuators. One of the key objectives was to derive expressions for the voltage and magnetic flux at which these devices switch. To do this, we have leveraged two variants of a state-of-the-art hybrid dynamical model, including magnetic saturation in one of them for enhanced accuracy at the cost of greater complexity. Both models assume a one-to-one relationship between current and magnetic flux. This simplification yields valuable analytical expressions for initial understanding, rapid analysis, and design optimization. For applications involving high-coercivity materials or high-frequency scenarios, a more comprehensive analysis incorporating magnetic hysteresis may be beneficial, although this would probably imply numerical procedures.